OpenStudy (anonymous):

What is the gcf for 18 and 84

OpenStudy (matlee):

The method is to figure out the prime factorization: 18 / 2 = 9 9 / 3 = 3 3 / 3 = 1 2 x 3 x 3 = 18 42 / 2 = 21 21 / 3 = 7 7 / 7 = 1 2 x 3 x 7 = 42 84 / 2 = 42 42 / 2 = 21 21 / 3 = 7 7 / 7 = 1 2 x 2 x 3 x 7 = 84 Look at each factorization and figure out which numbers appear in all the factorizations: 2 x 3 x 3 = 18 2 x 3 x 7 = 42 2 x 2 x 3 x 7 = 84 The common factors are 2 and 3: 2 x 3 = 6 Answer: GCF(18, 42, 84) = 6 P.S. Double-check that all are divisible by 6: 18 / 6 = 3 42 / 6 = 7 84 / 6 = 14

OpenStudy (matlee):

This is not m work